OpenStudy (anonymous):

simplify the equation: q/45-5/27=2/5

jhonyy9 (jhonyy9):

q/45 =2/5 +5/27 q/45 = (2*27+5*5)/135 q/45 = 79/135 q=79*45/135 =79*5/15 =79/3 q=79/3